I have an algebra question. On my worksheet, it says "Determine a relationship between the x- and y-values. Write an equation." I know how to do that, but for the second problem, the ordered pairs are (2,3), (3,5), (4,7), and (5,9). I put them in a table and it looks like the relationship is a sequence, because as you go up, you add 1 to 2 to make 3, and then you add 2 to 3 to make 5, and so on. How would you write this as an equation? (i.e. y=x+3)
This is what the problem looks like:
x 2 3 4 5
y 3 5 7 9

Found 2 solutions by AnlytcPhil, Sarpi:
Answer by AnlytcPhil(1806)

x 2 3 4 5
y 3 5 7 9

There are more than one way to do this kind of problem.
----------------------------------------
First way:

Look at (2,3).  If you add 1 to 3 you get 4, and 4 is twice 2.

Look at (3,5).  If you add 1 to 5 you get 6, and 6 is twice 3.

Look at (4,7).  If you add 1 to 7 you get 8, and 8 is twice 4.

Look at (5,9).  If you add 1 to 9 you get 10, and 10 is twice 5.

Now take the general case (x,y) and say the same thing about it:

Look at (x,y).  If you add 1 to y you get y+1, and y+1 is twice x.

Write that as an equation:

             y+1 = 2x

which simplifies to

               y = 2x - 1
------------------------------------------------------------

Second way:

Look at the point (2,3).  If you double 2 you get 4, and 4 is 1 more than 3.

Look at the point (3,5).  If you double 3 you get 6, and 6 is 1 more than 5.

Look at the point (4,7).  If you double 4 you get 8, and 8 is 1 more than 7.

Look at the point (5,9).  If you double 5 you get 10, and 10 is 1 more than 9.

Now take the general case (x,y) and say the same thing about it:

Look at the point (x,y).  If you double x you get 2x, and 2x is 1 more than y.

Write that as an equation:

                 2x = y + 1

which simplifies to

             2x - 1 = y                    

or 

                  y = 2x - 1

-------------------------------------------------------------------
Third way:

Look at the point (2,3).  If you subtract 2 from 3 you get 1, and 1 is 1 less than 2.

Look at the point (3,5).  If you subtract 3 from 5 you get 2, and 2 is 1 less than 3.

Look at the point (4,7).  If you subtract 4 from 7 you get 3, and 3 is 1 less than 4.

Look at the point (5,9).  If you subtract 5 from 9 you get 4, and 4 is 1 less than 5.

Now take the general case (x,y) and say the same thing about it:

Look at the point (x,y).  If you subtract x from y you get y-x, and y-x is 1 less than x.

Write that as an equation:

              y-x = x-1

which simplifies to

                y = 2x - 1
------------------------------------

There are even other ways to do it but you always get y = 2x - 1



Edwin

Answer by Sarpi(32)

X Y
2 3
3 5
4 7
5 9

From the data, X moves in steps of 1 point and Y in steps of 2 points. Therefore, if X is increased by 1, Y is increased by 2.

=> The figures are in sequence, so we need to determine how to go from X to Y, that is 2 to 3, 3 to 5, 4 to 7, and 5 to 9.
From this, it can be seen that the changes between X and Y are 1, 2, 3, and 4 respectively.

Interpretation;
from X value of 2 to Y value of 3 = 2+1
from X value of 3 to Y value of 5 = 3+2
from X value of 4 to Y value of 7 = 4+3
from X value of 5 to Y value of 9 = 5+4 ; where these adding numbers are the value of the previous X
(2+prev. X value = 3, 3+prev X value = 5 and so on)

Mathematically, the relationship is Y = Current value of X + Previous value of X
Y = X + (X - 1)
Y = X + X -1
Y = 2X - 1

Verification:
if X = 2 then Y = 2*2 - 1 = 3
if X = 3 then Y = 2*3 - 1 = 5
if X = 4 then Y = 2*4 - 1 = 7
if X = 5 then Y = 2*5 - 1 = 9